OpenStudy (anonymous):

Please assist me? I'll answer anything related to English, Psychology, or Philosophy in return? ;) Solve 2x + 5y = 10 for y. a.) y = 2x + 2 b.) y = -2x + 10 c.) y = x + 2 d.) y = x + 10

OpenStudy (anonymous):

(I think it's "C")

OpenStudy (anonymous):

Wait, let me revise, sorry!

OpenStudy (kropot72):

2x + 5y = 10 Subtracting 2x from both sides: 2x - 2x + 5y = 10 - 2x 5y = 10 - 2x Divide both sides by 5 to get: y = 2 - 2/5(x)

OpenStudy (kropot72):

\[y=2 -\frac{2x}{5}\]